OpenStudy (anonymous):

simplify: \[\left| 5-11 \right|\]

jimthompson5910 (jim_thompson5910):

5 - 11 is what

OpenStudy (anonymous):

what do u mean?

jimthompson5910 (jim_thompson5910):

what is 5-11 equal to

OpenStudy (anonymous):

-6

jimthompson5910 (jim_thompson5910):

so |5-11| = |-6|

jimthompson5910 (jim_thompson5910):

to evaluate |-6|, you just report the positive version of that number so |-6| = 6

jimthompson5910 (jim_thompson5910):

this means |5-11| = 6

jimthompson5910 (jim_thompson5910):

basically what |5-11| means is that 5 is 6 units away from 11 or 11 is 6 units away from 5 so that's why |5-11| = 6